OpenStudy (anonymous):

Calculate the wavelength of the radiation released when an electron moves from n=5 to n=2

OpenStudy (anonymous):

Is this for a Hydrogen atom?

OpenStudy (anonymous):

If so, the energy of a hydrogen energy level is:\[E(n)=\frac{-13.6}{n^2}eV\]So \[E(1)=\frac{-13.6}{1^2}=-13.6eV\]\[E(2)=\frac{-13.6}{2^2}=-3.4eV\]\[...\]\[E(5)=\frac{-13.6}{5^2}=-0.54eV\]The energy of the radiation is given by the difference between these energy levels, from which you can work out the wavelength. \[E_{Joules}=\frac{E_{eV}}{1.6\times 10^{-19}}\]\[\lambda=\frac{hc}{E_{Joules}}\]

OpenStudy (anonymous):

see that was the problem i was having the problem doesn't specify whether its for a hydrogen atom or not the professor ended up not counting the problem toward the grade but thank you!